A 78 kg man stands on a bathroom scale at the center of an elevator. As the elevator accelerates from rest downwards to -11.4 m/s in 3.04 s, what is the scale reading in Newtons?
Answer #1

Vi = initial velocity = 0 m/s

Vf = final velocity = - 11.4 m/s

t = time = 3.04 sec

acceleration is given as ::

a = (Vf - Vi) / t

a = -11.4 - 0 /3.04

a = -3.75 m/s2 in downward direction

The force equation for the motion of man in the elevator is given as

mg - Fn = ma

78(9.8) - Fn = 78(3.75)

Fn = 471.9 N